OpenStudy (anonymous):

Find the inverse transform of s/[(s^2+a^2)(s^2+b^2)], where a^2 is not equal to b^2 and ab does not equal 0. Please show two different ways to find the inverse transform.

ganeshie8 (ganeshie8):

have you tried partial fractions ?

OpenStudy (anonymous):

No I haven't but I'm sure that is one of the ways to solve it. I fear it will get very tricky very quickly lol

ganeshie8 (ganeshie8):

not really

OpenStudy (anonymous):

if partial fractions is one way....could convolution be another way?

ganeshie8 (ganeshie8):

\[\large \dfrac{s}{(s^2+a^2)(s^2+b^2)} = \dfrac{s(s^2+b^2-s^2-a^2)}{(s^2+a^2)(s^2+b^2)}\]

ganeshie8 (ganeshie8):

that simplifies nicely but its not correct, it needs some more tweaking

ganeshie8 (ganeshie8):

looks there is an extra (b^2-a^2) in the numerator ?

OpenStudy (anonymous):

yeah I was just wondering about that lol

ganeshie8 (ganeshie8):

can we say \[\large \dfrac{s}{(s^2+a^2)(s^2+b^2)} = \dfrac{1}{b^2-a^2}\dfrac{s(s^2+b^2-s^2-a^2)}{(s^2+a^2)(s^2+b^2)}\]

ganeshie8 (ganeshie8):

\[\begin{align} \dfrac{s}{(s^2+a^2)(s^2+b^2)} &= \dfrac{1}{b^2-a^2}\dfrac{s(s^2+b^2-s^2-a^2)}{(s^2+a^2)(s^2+b^2)}\\~\\ &= \dfrac{1}{b^2-a^2}\left[\dfrac{s}{s^2+a^2}-\dfrac{s}{s^2+b^2}\right]\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

does that look okay

OpenStudy (anonymous):

that looks correct from the step above it if that's what you mean. so from there would it be....something times cos(at)-cos(bt)...?

ganeshie8 (ganeshie8):

Yep! thats it

OpenStudy (anonymous):

so my final answer is (1/b^2-a^2)*(cos(at)-cos(bt))? I leave it like that?

OpenStudy (anonymous):

will I get the same answer by convolution....I have to be able to solve it two ways for my class lol

OpenStudy (sidsiddhartha):

yes convolution will produce same result then u have to take \[F_1(s)=\frac{ s }{ s^2+b^2}\\F_2(s)=\frac{ 1 }{ s^2+a^2 }\\then ~simply\\ \int\limits_{0}^{t}f_1(u)*f_2(t-u)du\]

OpenStudy (sidsiddhartha):

\[where\\f_1(u)=L^{-1}F_1(t)|_{s=u}\\f_2(u)=L^{-1}F_2(t)|_{s=t-u}\]

OpenStudy (anonymous):

awesome thanks!

OpenStudy (anonymous):

@sidsiddhartha I am a bit confused on the convolution part. How would I take the inverse laplace of 1/(s^2+a^2) so I could do the rest of the problem?

OpenStudy (sidsiddhartha):

ok we have\[F_2(s)=\frac{ 1 }{ s^2+a^2 }\\so\\f_2(t)=L^{-1}[F_2(s)]=L^{-1}[\frac{ 1 }{ a }*\frac{ a }{ a^2+s^2 }]\\=\frac{ 1 }{ a }*L^{-1}\frac{ a }{ a^2+s^2 }\] make sense?

OpenStudy (anonymous):

I think so. So would the inverse be 1/a*cos(at)?

OpenStudy (sidsiddhartha):

sin

OpenStudy (anonymous):

no wait that would be sin instead of cos I think

OpenStudy (sidsiddhartha):

yup right :)

OpenStudy (anonymous):

ok! So from here I just turn the t's into u's and put them into the formula to solve the convolution?

OpenStudy (sidsiddhartha):

yes and \[f_1(t)=\cos(bt)\] then turn the t's into u's and put them into the formula to solve the convolution as u said :)

OpenStudy (anonymous):

I know how to do the convolution when finding the laplace transform....is it the same process for finding the inverse laplace transform?

OpenStudy (sidsiddhartha):

yes basically tha same thing

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

ok sorry one more question...the formula becomes integral from 0 to t cos(bu)*(1/a)sin(t-au)du? that last part is throwing me off